OpenStudy (anonymous):

A constant force F=8i+5j−7k moves an object along a straight line from point (3, -6, -6) to point (-4, -8, 9). Find the work done if the distance is measured in meters and the magnitude of the force is measured in newtons.

OpenStudy (anonymous):

@e.mccormick can you take a look?

OpenStudy (e.mccormick):

OK. So, you know your distance formula for this?

OpenStudy (anonymous):

Well I know w = F • D So what I have done is found that f = (8,5,-7) and D = (-4, -8, 9) - (3, -6, -6) = (-7,-2,15) so I took (-7,-2,15) • (8,5,-7) = -59 but thats seems to be wrong

OpenStudy (e.mccormick):

Ah, you did d wrong. \(d(p,q)=\sqrt{(p_1-q_1)^2+(p_2-q_2)^2+\ldots+(p_n-q_n)^2}\)

OpenStudy (e.mccormick):

Then again... I didn't do physics, so if you got the other from a formula. I learned vectors through linear algebra. Hehe.

OpenStudy (anonymous):

Well this isnt physics just calc 3. Ill give it a shot

OpenStudy (anonymous):

I was looking at this: http://answers.yahoo.com/question/index?qid=20110907173249AA6wkZV

OpenStudy (e.mccormick):

Ah, may be a good shortcut: http://en.wikipedia.org/wiki/Dot_product

OpenStudy (e.mccormick):

"Mechanical work is the dot product of force and displacement vectors."

OpenStudy (e.mccormick):

So displacement rather than distance.

OpenStudy (e.mccormick):

And that would be the right one then.

OpenStudy (anonymous):

they way i did it would be the right one?

OpenStudy (e.mccormick):

Yes, because you found the displacment and they gave you force. As long as you did the do produc correctly, that would work.

OpenStudy (anonymous):

I did the math wrong. its -171

OpenStudy (e.mccormick):

Yah. I just found that myself.

OpenStudy (anonymous):

Thanks again

OpenStudy (e.mccormick):

np. Dot product has some really interesting uses.